Methanol(CH3OH)can be manufactured by combining gaseous carbon monoxide and hydrogen gas.If I combine 9.13kg of carbon monoxide with 12.5kg of hydrogen gas show the balanced equation,find the limiting reagent ,and calculate how much product I’ll create.Finally show much unused reactant I will have left over.
1
Expert's answer
2020-03-16T12:14:49-0400

CO + 2H2 -> CH3OH

CO is limiting reagent

n (CO) = m/M = 9.13 kg / 28.01 g/mol = 0.33 kmol

n(CH3OH) = n(CO)

m (CH3OH) = nM = 0.33 kmol * 32.04 = 10.44 kg

n (H2) = 2n (CO) = 0.66 kmol

m (H2) used = nM = 0.66 kmol * 1 g/mol = 0.66 kg

m (H2) unused = 12.5 kg - 0.66 kg = 11.84 kg